1) Sue is both the 50th best & the 50th worst student at her school. How many
students attend her school ?
2) In a race from point X to point Y & back, Jack averages 30 miles per hour to
point Y & 10 miles per hour back to point X. Sandy averages 20 miles per
hour in both directions. Between Jack & Sandy, who finished first ?

Answer by usyim88hk(157)

1) 99. because 49 people are better then her and 49 people are worse than her and so 49+49+1(Sue) = 99
2)They both finish at the same time because 30+10 = 20+20